# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

### v1

@@ A.2 An Alternative Bound for the Convergence Rate of the Dynamic Game
In this appendix, we provide an alternative bound for the convergence rate of Algorithm 1 in the dynamic load-balancing game, which, for certain ranges of game parameters, could be better than the convergence rate given in Theorem 4.11.
@@
As before, the main idea is to analyze how quickly $\sum_{j=1}^{m}s_{j}^{t}$ approaches zero. Define $\lambda_{\text{min}}=\min_{i}\lambda_{i}$ and $\lambda_{\text{max}}=\max_{i}\lambda_{i}$. Similarly, let $\mu_{\text{min}}=\min_{j}\mu_{j}$ and $\mu_{\text{max}}=\max_{j}\mu_{j}$. To lower-bound the amount of decrease in $\big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\big{|}$, we consider two cases:
@@
      $s_{m}^{t}<\mu_{\text{min}}$: In this case we have $s_{m}^{t+1}=0$, thus $s_{m}^{t+1}-s_{m}^{t}=-s_{m}^{t}$. Moreover, since server $m$ has not received any job at time $t$, according to (13), we have $\frac{s_{m}^{t}}{\mu_{\text{min}}}\geq\frac{\lambda_{\text{min}}}{\sum_{j=1}^{m-1}\mu_{j}}$. This gives us
      $\frac{\mu_{\text{min}}\lambda_{\text{min}}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\text{min}}}\leq\Big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\Big{|}.$
  • All servers are receiving jobs, but the state has not yet converged to zero: According to Theorem 4.11, in this case we have
@@
    $\sum_{j=1}^{m}\mu_{j}-\lambda_{\text{max}}\leq\Big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\Big{|}.$
@@
$\leq\frac{\sum_{j=1}^{m}s_{j}^{0}}{\min\big{\{}\sum_{j=1}^{m}\mu_{j}-\lambda_{\max},\ \mu_{\min},\ \frac{\mu_{\text{min}}\lambda_{\min}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\min}}\big{\}}}.$

### v2

@@ A.3 An Alternative Bound for the Convergence Rate of the Dynamic Game
In this appendix, we provide an alternative bound for the convergence rate of Algorithm 1 in the dynamic load-balancing game, which, for certain ranges of game parameters, could be better than the convergence rate given in Theorem 11.
@@
As before, the main idea is to analyze how quickly $\sum_{j=1}^{m}s_{j}^{t}$ approaches zero. Define $\lambda_{\text{min}}=\min_{i}\lambda_{i}$ and $\lambda_{\text{max}}=\max_{i}\lambda_{i}$. Similarly, let $\mu_{\text{min}}=\min_{j}\mu_{j}$ and $\mu_{\text{max}}=\max_{j}\mu_{j}$. To lower-bound the amount of decrease in $\left|\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\right|$, we consider two cases:
@@
      $s_{m}^{t}<\mu_{\text{min}}$: In this case we have $s_{m}^{t+1}=0$, thus $s_{m}^{t+1}-s_{m}^{t}=-s_{m}^{t}$. Moreover, since server $m$ has not received any job at time $t$, according to Eq. (14), we have $\frac{s_{m}^{t}}{\mu_{\text{min}}}\geq\frac{\lambda_{\text{min}}}{\sum_{j=1}^{m-1}\mu_{j}}$. This gives us
      $\frac{\mu_{\text{min}}\lambda_{\text{min}}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\text{min}}}\leq\left|\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\right|.$
  • All servers are receiving jobs, but the state has not yet converged to zero: According to Theorem 11, in this case we have
@@
    $\sum_{j=1}^{m}\mu_{j}-\lambda_{\text{max}}\leq\left|\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\right|.$
@@
$\leq\frac{\sum_{j=1}^{m}s_{j}^{0}}{\min\left\{\sum_{j=1}^{m}\mu_{j}-\lambda_{\max},\ \mu_{\min},\ \frac{\mu_{\text{min}}\lambda_{\min}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\min}}\right\}}.$